N 10.00 polnts PA 15-6 La Villa is a ski resort in the Italian Alps... La Villa is a ski resort in the Italian Alps. Many guests visit the town, however, most of them come in the winter season because of the phenom initial forecast of 17,000, forecast demand for the first four quarters in 2015. (Round Seasonal Index to 3 decimal places and all the other answers to 2 decimal places.) Deaseasonalized Smoothed Demand Re-seasonalized Forecast Seasonal Year Qtr Demand Quarter Season Average Forecast Index Demand 17,720.43 17,000.00 17,812.27 2012 1 22,100 1 21,866.67 1.247 2015Q1 2 14,900 2 14,666.67 0.837 18,275.00 2015Q2 3 13,700 3 14,466.67 0.825 16,604.15 17,431.25 2015Q3 4 19,000 4 19,133.33 1.091 17,411.15 16,498.44 2015Q4 2013 1 20,100 Overall 17,533.33 16,118.68 17,123.83 2 17,000 20,310.63 17,867.87 3 16,500 20,000.00 17,650.90 4 18,700 17,140.24 17,363.18 2014 1 23,400 18,765.04 17,697.38 14,456.39 19,123.04 16,000.00 18,056.83 2 12,100 3 13,200 17,367.28 4 19,700 16,325.46
(of 10) Save & Exit Submit eason because of the phenomenal sking. A group of hotels has analyzed their demand for the last three years. Fill in the table below. Using a smoothing parameter of 0.25 and a Re-seasonalized t Forecast
Answer #1

Year

Otr

Demand

Quarter

Season Average

Seasonal Index

Deseasonalized Demand

Smoothed Demand

Forecast

Quarter

Re-Seasonalized Forecast

2012

1

22100

1

21866.66667

1.247148289

17720.43

17000

2015

1

17423.6115

2012

2

14900

2

14666.66667

0.836501901

17812.27

18275

2015

2

17433.223

2012

3

13700

3

14466.66667

0.825095057

16604.15

17431.25

2015

3

17442.8345

2012

4

19000

4

19133.33333

1.091254753

17411.15

16498.44

2015

4

17452.446

2013

1

20100

Overall

17533.33333

16116.77

17123.83

2013

2

17000

20322.73

17867.87

2013

3

16500

19997.70

17650.9

2013

4

18700

17136.24

17363.18

2014

1

23400

18762.80

17697.38

2014

2

12100

14465.00

19123.04

2014

3

13200

15998.16

17367.28

2014

4

19700

18052.61

16325.46

Excel Formula Sheet: Copy & Paste in Cell A1

Year

Otr

Demand

Quarter

Season Average

Seasonal Index

Deseasonalized Demand

Smoothed Demand

Forecast

Quarter

Re-Seasonalized Forecast

2012

1

22100

1

=AVERAGE(C2,C6,C10)

=F2/$F$6

=C2/G2

17000

2015

1

=9.6115*K2 + 17414

2012

2

14900

2

=AVERAGE(C3,C7,C11)

=F3/$F$6

=C3/G3

18275

2015

2

=9.6115*K3 + 17414

2012

3

13700

3

=AVERAGE(C4,C8,C12)

=F4/$F$6

=C4/G4

17431.25

2015

3

=9.6115*K4 + 17414

2012

4

19000

4

=AVERAGE(C5,C9,C13)

=F5/$F$6

=C5/G5

16498.44

2015

4

=9.6115*K5 + 17414

2013

1

20100

Overall

=AVERAGE(F2:F5)

=C6/G2

17123.83

2013

2

17000

=C7/G3

17867.87

2013

3

16500

=C8/G4

17650.9

2013

4

18700

=C9/G5

17363.18

2014

1

23400

=C10/G2

17697.38

2014

2

12100

=C11/G3

19123.04

2014

3

13200

=C12/G4

17367.28

2014

4

19700

=C13/G5

16325.46


y 9.6115x+17414 Chart Title R2 0.0021 19500 19000 18500 18000 17500 17000 16500 16000 15500 15000 14500 1 2 3 4 7 9 10 11 12

C A D G H Demand Quarter Season Average Seasonal Index Deseasonalized Demand Smoothed Demand Forecast Quarter Re-Seasonalized